# Turán number of complete multipartite graphs in multipartite graphs

### v1

@@ 3. $(K3)$
Suppose to the contrary, there is a vertex $v$ in $\tilde{U}_{1}\setminus(Z_{1}^{1}\cup W_{1}^{\prime})$ or $W_{1}^{\prime}\setminus V_{q_{1}}$, namely, $v\in Z_{i}^{1}$ for some $2\leq i\leq r$ or $v\in W_{1}^{\prime}\setminus V_{q_{1}}$. Suppose $v\in V_{l}$. Then $l\neq q_{1}$. Morepver, if $i$ is defined, then $W^{\prime}_{1}\cap V_{q_{1}}\subseteq V\setminus(\tilde{Z}_{i}\cup V_{l})$; otherwise, $W^{\prime}_{1}\cap V_{q_{1}}\subseteq V\setminus V_{l}$. By (Deg), we have $\overline{d}(v,W_{1}^{\prime}\cap V_{q_{1}})\leq 2\varepsilon n$. Let $N:=W_{1}^{\prime}\cap V_{q_{1}}\cap N(v)$. We have $|(W_{1}^{\prime}\cap V_{q_{1}})\setminus N|\leq 2\varepsilon N$. Since $|W_{1}^{\prime}\setminus V_{q_{1}}|\leq|W_{1}^{\prime}\setminus W_{1}|\leq 2\gamma n$, it follows that $|W^{\prime}_{1}\setminus N|\leq 2\varepsilon n+2\gamma n\leq 3\varepsilon n$.
Recall (4.15), $|W_{1}|>3t\varepsilon n$. By $(K3)$ (if $v\in\tilde{Z}_{1}\setminus Z_{1}$) or $(K1)$ (if $v\in W_{1}^{\prime}\setminus V_{q_{1}}$), we know that $G[Z_{1}^{1},N]$ contains no $K_{t-1,t}$ with the part of size $t$ in $N$. This implies that $e_{G}(Z_{1}^{1},N)=O(n^{2-1/(t-1)})$. Furthermore, by $(P3)$ and $(K1)$, $G[\tilde{Z}_{1}\setminus Z_{1}^{1},Z_{1}^{1}]$ is a $K_{t,t}$-free graph with one part of size at most $(r-1)\sqrt{\gamma}n$ and the other part of size at most $an$. Thus, $e_{G}(\tilde{Z}_{1}\setminus Z_{1}^{1},Z_{1}^{1})\leq C(r-1)\sqrt{\gamma}n(an)^{1-1/t}$. By the similar arguments, we have $e_{G}(W_{1}^{\prime}\setminus N,Z_{1}^{1})\leq C(3\varepsilon n)(an)^{1-1/t}$.
Putting these bounds together with $e_{G}(Z_{1}^{1})\leq z_{t}^{(a)}(n)$, we get
$f_{1}$ $=e_{G}(Z_{1}^{1})+e_{G}(Z_{1}^{1},N)+e_{G}(\tilde{Z}_{1}\setminus Z_{1}^{1},Z_{1}^{1})+e_{G}(W_{1}^{\prime}\setminus N,Z_{1}^{1})$
$\leq z_{t}^{(a)}(n)+O(n^{2-1/(t-1)})+O(\sqrt{\gamma}n^{2-1/t})+O(\varepsilon n^{2-1/t}).$
By ($E1$), this implies that $f_{1}<z_{t}^{(a+1)}(n)-\varepsilon n^{2-1/t}$, contradicting (4.14). ∎
@@
$\alpha_{1}=\beta_{1}=0,\quad\text{and}\quad e_{G}(\tilde{U}_{1})=f_{1}\leq z_{t}^{(a+1)}(|Z_{1}^{1}\cap V_{1}|,\dots,|Z_{1}^{1}\cap V_{a}|,|W^{\prime}_{1}|)\leq z_{t}^{(a+1)}(n).$ (4.16)
@@
    If there is a copy of $K_{1,t}$ in $\tilde{U}_{i}\setminus(Z_{1}\cup V_{q_{1}})$, then there exists $l\in[r]\setminus\{i\}$ such that the vertex set of $K_{1,t}$ intersects $V_{q_{l}}$ and every cluster in $Z_{l}$.
@@
    $\tilde{Z}_{i}\setminus Z_{1}$ and $Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})$ are $K_{1,t}$-free.
  3. (3)
    If $t<a$, then $\tilde{U}_{i}\setminus(Z_{1}\cup V_{q_{1}})$ is $K_{1,t}$-free.
@@
First assume that $N$ is $K_{t-1,t}$-free and thus $e_{G}(N)=O(n^{2-1/(t-1)})$. Note that, since $|\tilde{U}_{1}\setminus N|\leq(t+1)2\varepsilon n$, the edges in $\tilde{U}_{1}$ incident to $\tilde{U}_{1}\setminus N$ can be split into $a+1$ bipartite $K_{t,t}$-free graphs each with one part of size at most $(t+1)2\varepsilon n$ and the other part of size at most $an$. Thus, the number of such edges is $O(\varepsilon n^{2-1/t})$. This gives $f_{1}=O(n^{2-1/(t-1)})+O(\varepsilon n^{2-1/t})$, contradicting (4.14).
@@
We thus assume $N$ contains a copy of $K_{t-1,t}$. Together with $B$, they form a copy of $K_{3}{(t)}$ in $G[\tilde{U}_{1}\cup\tilde{U}_{i}]$ and we denote its vertex set by $B^{\prime}$. By Claim 4.6, there exists $l\notin\{1,i\}$ such that $B^{\prime}$ intersects $V_{q_{l}}$ and every cluster of $Z_{l}$. By Claim 4.8, $\tilde{U}_{1}\cap U_{l}=\emptyset$, so $B^{\prime}\cap Z_{l}=B\cap Z_{l}$ and $B$ indeed intersects every cluster of $Z_{l}$. Since $\tilde{U}_{i}\cap Z_{l}\supseteq B\cap Z_{l}\neq\emptyset$, we infer that $|W_{l}|\geq(1-2\varepsilon)n$ from $(Q3)$, which implies that $q_{l}\neq q_{1}$ because of (4.15). It follows that $W_{1}\cap V_{q_{l}}=\emptyset$ and thus $B\cap V_{q_{l}}=B^{\prime}\cap V_{q_{l}}\neq\emptyset$, as desired.
@@
For Part (2), let $A_{i}:=Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})$ and $B$ be the vertex set of a copy of $K_{1,t}$ in $\tilde{Z}_{i}\setminus Z_{1}$ or in $A_{i}$. Then, by the first part of the claim, there exists $l\in[r]\setminus\{i\}$ such that $B$ intersects $V_{q_{l}}$ and every cluster in $Z_{l}$. This is impossible if $B\subseteq A_{i}$ because $A_{i}\cap Z_{l}=\emptyset$ for any $l\notin\{1,i\}$, and also impossible if $B\subseteq\tilde{Z}_{i}\setminus Z_{1}$ because in which case $B\cap W=\emptyset$ and thus $B\cap V_{q_{l}}=\emptyset$ for any $l\notin\{1,i\}$.
Part (3) follows from Part (1) immediately. ∎
The next claim bounds $e_{G}(Z_{1}^{i})$ for $i>1$.
Claim 4.10.
For every $i\in[2,r]$, $e_{G}(Z_{1}^{i})\leq\binom{a}{2}(t-1)|Z_{1}^{i}|$.
Proof.
Suppose $|W_{1}|\geq(1-\varepsilon-\sqrt{\gamma})n$ (otherwise $Z_{1}^{i}=\emptyset$ for $i>1$ by $(Q2)$ and there is nothing to prove). Fix $i\in[2,r]$. Suppose $a\geq 2$ (if $a=1$ then $e_{G}(Z_{1}^{i})=0$). We claim that for distinct $i_{1},i_{2}\in[a]$, there can not be two copies of $K_{1,t}$ in the bipartite graph $G[Z_{1}^{i}\cap V_{i_{1}},Z_{1}^{i}\cap V_{i_{2}}]$, one centered in $Z_{1}^{i}\cap V_{i_{1}}$ and the other centered in $Z_{1}^{i}\cap V_{i_{2}}$. We show this for $i_{1}=1$ and $i_{2}=2$. Suppose $w\in Z_{1}^{i}\cap V_{1}$ and $u_{1},\dots,u_{t}\in Z_{1}^{i}\cap V_{2}$ form a copy of $K_{1,t}$ while $w^{\prime}\in Z_{1}^{i}\cap V_{2}$, $u_{1}^{\prime},\dots,u_{t}^{\prime}\in Z_{1}^{i}\cap V_{1}$ form the other copy of $K_{1,t}$. By (Deg), $w,u_{1}^{\prime},\dots,u_{t}^{\prime}$ each has at most $2\varepsilon n$ non-neighbors in $\tilde{U}_{1}\setminus V_{1}$ and $w^{\prime},u_{1},\dots,u_{t}$ each has at most $2\varepsilon n$ non-neighbors in $\tilde{U}_{1}\setminus V_{2}$. Let
$N_{1}=\left(\tilde{U}_{1}\setminus(V_{1}\cup V_{2})\right)\cap N(w)\cap\bigcap_{i\in[t]}N(u_{i})\quad\text{ and }\quad N_{2}=\left(\tilde{U}_{1}\setminus V_{2}\right)\cap\bigcap_{i\in[t]}N(u_{i})$
(thus $N_{1}\cap N_{2}\subseteq\tilde{U}_{1}$). By $(P3)$ and $|W_{1}\cap W_{1}^{\prime}|\geq(1-\varepsilon-\sqrt{\gamma})n-2r\gamma n\geq(1-2\varepsilon)n$, we have
$|N_{1}|\geq(1-\sqrt{\gamma})an-2n+(1-2\varepsilon)n-(t+1)2\varepsilon n\geq(a-1)n-(2t+5)\varepsilon n$
and similarly $|N_{2}|\geq an-(2t+3)\varepsilon n$. If $G[N_{1},N_{2}]$ contains a copy of $K_{t-1,t}$ with the part of size $t$ in $N_{1}$, then it together with $\{w,u_{1},\dots,u_{t}\}$ forms a copy of $K_{3}{(t)}$ in $\tilde{U}_{1}\cup\tilde{U}_{i}$. Let $B$ be the vertex set of this $K_{3}{(t)}$ and note that $B\subseteq U_{1}$. Thus $B\cap Z_{l}=\emptyset$ for $l\neq 1$ contradicting Claim 4.6. This implies that $e_{G}(N_{1},N_{2})=O(n^{2-1/(t-1)})$. Now let $N_{1}^{\prime}=(\tilde{U}_{1}\setminus(V_{1}\cup V_{2}))\cap N(w^{\prime})\cap\bigcap_{i\in[t]}N(u_{i}^{\prime})$ and $N_{2}^{\prime}=(\tilde{U}_{1}\cap V_{2})\cap\bigcap_{i\in[t]}N(u_{i}^{\prime})$. By the same argument as above, $G[N_{1}^{\prime},N_{2}^{\prime}]$ contains no copy of $K_{t-1,t}$ with the part of size $t$ in $N_{1}^{\prime}$, which implies that $e_{G}(N_{1}^{\prime},N_{2}^{\prime})=O(n^{2-1/(t-1)})$. Note that all but $O(\varepsilon n^{2-1/t})$ edges of $G[\tilde{U}_{1}]\setminus G[V_{1},V_{2}]$ are in $G[N_{1}^{\prime},N_{2}^{\prime}]$ or $G[N_{1},N_{2}]$. We thus derive that $f_{1}\leq z_{t}^{(2)}(n)+O(n^{2-1/(t-1)})+O(\varepsilon n^{2-1/t})<z_{t}(n)-\varepsilon n^{2-1/t}$ by ($E1$), contradicting (4.14).
The conclusion above implies that $e_{G}[Z_{1}^{i}\cap V_{i_{1}},Z_{1}^{i}\cap V_{i_{2}}]\leq(t-1)\max\{|Z_{1}^{i}\cap V_{i_{1}}|,|Z_{1}^{i}\cap V_{i_{2}}|\}$. Therefore,
$e_{G}(Z_{1}^{i})=\sum_{i_{1},i_{2}\in[a]}e_{G}[Z_{1}^{i}\cap V_{i_{1}},Z_{1}^{i}\cap V_{i_{2}}]\leq\sum_{i_{1},i_{2}\in[a]}(t-1)|Z_{1}^{i}|\leq\binom{a}{2}(t-1)|Z_{1}^{i}|.\qed$
Claims 4.9 (2) and 4.10 together give $e_{G}(Z_{j}^{i})\leq\max\{(t-1)|Z_{j}^{i}|,\binom{a}{2}(t-1)|Z_{j}^{i}|\}=\binom{a}{2}(t-1)|Z_{j}^{i}|$ whenever $i\neq j$. Thus, by $(P3)$,
$\sum_{i,j:i\neq j}e_{G}(Z_{j}^{i})\leq r(r-1)\binom{a}{2}(t-1)\sqrt{\gamma}n\leq{a}^{2}r^{2}t\sqrt{\gamma}n$ (4.17)
Together with $f_{1}\leq z^{(a+1)}_{t}(n)$ and $f_{i}\leq(t-1)|\tilde{U}_{i}|$ for $i\geq 2$, we derive from (4.9) that
$e(G)=t_{r}(k)n^{2}+z^{(a+1)}_{t}(n)+O(n).$
This concludes the proof of Theorem 5 (note that we have not used  ($E2$) or ($E3$)).
Proof of Theorem 6
We refine our earlier estimates and prove Theorem 6. Since $\sum_{i\neq j}e_{G}(Z_{j}^{i})\leq a^{2}r^{2}t\sqrt{\gamma}n$ and $f_{j}\leq(t-1)|\tilde{U}_{j}|$ for $j\in[2,r]$, we have the following bound better than (4.14),
$f_{1}\geq z_{t}(n)-3rC_{0}kn.$ (4.18)
Furthermore, we claim that
$\text{if}\quad b>0,\quad\text{then}\quad|W_{1}^{\prime}|\geq(1-\gamma)n\quad(\text{and thus}\quad 1\in L).$ (4.19)
Indeed, by Claim 4.8, if $|W_{1}^{\prime}|<(1-\gamma)n$, then we have $f_{1}\leq z_{t}^{(a+1)}(n,\dots,n,|W_{1}^{\prime}|)\leq z_{t}^{(a+1)}(n)-\delta n^{2-1/t}$ for some $\delta>0$ by ($E2$). This contradicts (4.18).
Next we show that $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$.
Claim 4.11.
Suppose $v_{0}\in V(G)$ and $i\in[r]$ satisfy that $v_{0}$ has at least $\varepsilon n$ neighbors in $Z_{j}$ for every $j\neq i$. Then $v_{0}$ has less than $\varepsilon n$ neighbors in $U_{i}$. In particular, we have $Z_{i}^{\prime\prime}=\emptyset$ for all $i\in[r]$ and $W^{\prime\prime}=\emptyset$.
Proof.
The second part of the claim follows immediately from the definitions of $Z_{i}^{\prime\prime}$ and $W^{\prime\prime}$.
Suppose to the contrary, that there exist $v_{0}\in V(G)$ and $i\in[r]$ such that $v_{0}$ has at least $\varepsilon n$ neighbors in $Z_{j}$ for every $j\neq i$ and at least $\varepsilon n$ neighbors in $U_{i}$. Since $|Z_{j}^{j}|\geq(1-\sqrt{\gamma})an$ for all $j\in[r]$, there exist sets $N_{1},\dots,N_{r-1}$ each of size $\varepsilon n-\sqrt{\gamma}an$ such that $N_{j}\subseteq Z_{j}^{j}\cap N(v_{0})$ for $j\neq i$ and $N_{i}\subseteq(Z_{i}^{i}\cup W_{i})\cap N(v_{0})$. Recall that $W_{1}^{\prime}=W_{1}^{\prime}\cap V_{q_{1}}$. By averaging, there exists $N_{1}^{\prime}\subseteq N_{1}$ with $|N_{1}^{\prime}|\geq(\varepsilon n-\sqrt{\gamma}an-2r\gamma n)/(a+1)\geq\varepsilon n/(a+2)$ such that all vertices of $N_{1}^{\prime}$ are in $Z_{1}^{1}\cup W_{1}^{\prime}$ and from the same cluster, that is,
$N_{1}^{\prime}\subseteq Q,\text{ where }Q\in\{V_{1}\cap Z_{1}^{1},V_{2}\cap Z_{1}^{1},\dots,V_{a}\cap Z_{1}^{1},W_{1}^{\prime}\}.$
Note that $N_{1}^{\prime}\subseteq W_{1}^{\prime}$ is possible only if $i=1$. If $i\neq 1$, then let $N_{i}^{\prime}:=N_{i}\setminus((W_{i}\setminus W_{i}^{\prime})\cup V_{q_{1}})$ and for every $j\in[r]\setminus\{1,i\}$, let $N_{j}^{\prime}:=N_{j}$. By $(P1)$, $|W_{i}\setminus W_{i}^{\prime}|\leq 2r\gamma n$, and by (4.19), $|W_{i}\cap V_{q_{1}}|\leq\gamma n$. Thus, we have $|N_{j}^{\prime}|\geq\varepsilon n/(a+2)$ for all $j\in[r]$. Because the sets $N_{j}^{\prime}$ are small, we can not apply the degree conditions (Deg) to them and instead, we use (D0).
Recall that $B_{1}$ is given by (D0). Next we show that $G[\tilde{U}_{1}\setminus B_{1}]$ does not contain a copy of $K_{t-1,t}$ such that the part of size $t$ is in $N_{1}^{\prime}$. Suppose instead, there is such a copy of $K_{t-1,t}$, with parts denoted by $A$ and $B$, such that $|A|=t$ and $A\subseteq N_{1}^{\prime}\setminus B_{1}$ and $B\subseteq\tilde{U}_{1}\setminus B_{1}$. Recall that $N_{i}^{\prime}\cap V_{q_{1}}=\emptyset$ and for each $j\in[r]\setminus\{1,i\}$, $N_{j}^{\prime}\subseteq Z_{j}^{j}$. Observe that for every $v\in\tilde{U}_{1}\setminus B_{1}$, we have $d(v,N_{j}^{\prime})\geq|N_{j}^{\prime}|-\sqrt{\gamma}n$. Indeed, if $j\neq i$, then $N_{j}^{\prime}\subseteq Z_{j}^{j}$ and we have $d(v,N_{j}^{\prime})\geq|N_{j}^{\prime}|-\sqrt{\gamma}n$ by (D0); otherwise note that $N_{i}^{\prime}\subseteq Z_{i}^{i}\cup(W_{i}^{\prime}\cap W_{i})$, and by (D0) and $N_{i}^{\prime}\cap V_{q_{1}}=\emptyset$ we have $d(v,N_{i}^{\prime})\geq|N_{i}^{\prime}|-\sqrt{\gamma}n$. Therefore, we obtain that the vertices in $A\cup B$ have at least $|N_{j}^{\prime}|-(2t-1)\sqrt{\gamma}n\geq(1-\gamma^{1/3})|N_{j}^{\prime}|$ common neighbors in each $N_{j}^{\prime}$, $j\in[2,r]$. Because each pair $N_{j}^{\prime},N_{j^{\prime}}^{\prime}$ has a high density, we can find a copy of $K_{r-1}{(t)}$ in the union of these common neighborhoods, which together with $A\cup B\cup\{v_{0}\}$ form a copy of $K_{r+1}(t)$, a contradiction.
Now given that $G[\tilde{U}_{1}\setminus B_{1}]$ does not contain a copy of $K_{t-1,t}$ such that the part of size $t$ is in $N_{1}^{\prime}\setminus B_{1}$, we shall give a refined estimate on $f_{1}$. Indeed, for each $V_{j}$, $j\in[a]$, we know that $G[N_{1}^{\prime}\setminus B_{1},(V_{j}\cap Z_{1}^{1})\setminus B_{1}]$ does not contain a copy of $K_{t-1,t}$ such that the part of size $t$ is in $N_{1}^{\prime}\setminus B_{1}$, implying that $e_{G}(N_{1}^{\prime}\setminus B_{1},(V_{j}\cap Z_{1}^{1})\setminus B_{1})=O(n^{2-1/(t-1)})$. Similarly we also have $e_{G}(N_{1}^{\prime}\setminus B_{1},W_{1}^{\prime}\setminus B_{1})=O(n^{2-1/(t-1)})$. Suppose $N_{1}^{\prime}\subseteq V_{q}$ for some $q\in[a]\cup\{q_{1}\}$, then we have
Recall that $|N_{1}^{\prime}|\geq\varepsilon n/(a+2)$ and $|B_{1}|\leq 2\sqrt{\gamma}n$. Therefore, we can bound $f_{1}\leq|E(G[\tilde{U}_{1}])|$ by
$f_{1}\leq z_{t}\left((1-\tfrac{\varepsilon}{a+2})n,n,\dots,n\right)+O(n^{2-1/(t-1)})+O(\sqrt{\gamma}n^{2-1/t})<z_{t}(n)-3rC_{0}kn,$
When $b=0$, since $Z^{\prime\prime}_{i}=\emptyset$ for all $i\in[r]$, we can improve  (4.13) to $e(G)\leq t_{r}(k)n^{2}+z_{t}^{(a)}(n)+(r-1)\lfloor\frac{t-1}{2}an\rfloor$, proving Theorem 6 for $b=0$.
In the rest of the proof, we assume $b>0$. We start with the following claim.
@@
Suppose instead, for some $i_{0}\in[2,r]$ with $|W_{i_{0}}|\geq 2\varepsilon n$, there exists $v\in\tilde{U}_{i_{0}}\setminus(U_{i_{0}}\cup V_{q_{i_{0}}})$. By $(P4)$ and the fact that $v\in\tilde{U}_{i_{0}}\setminus U_{i_{0}}$, we infer that $d(v,Z_{j})\geq\varepsilon n$ for all $j\neq{i_{0}}$. Then, by Claim 4.11, we have $d(v,U_{i_{0}})<\varepsilon n$. Consequently, $d(v,W_{i_{0}}^{\prime}\cap W_{i_{0}})<\varepsilon n$, namely, $v$ has at least $2\varepsilon n-2\gamma n-\varepsilon n\geq(3/4)\varepsilon n$ non-neighbors in $W_{i_{0}}^{\prime}\cap W_{i_{0}}$ (in $G$). Note that $v$ is adjacent to all the vertices of $W_{i_{0}}^{\prime}\cap W_{i_{0}}$ in $T$. Since $G[W^{\prime}_{i_{0}}]\subseteq T[W^{\prime}_{i_{0}}]$, we infer that
$e_{G}(W_{i_{0}}^{\prime})+e_{G}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}},W_{i_{0}}^{\prime})\leq e_{T}(W_{i_{0}}^{\prime})+|\tilde{Z}_{i_{0}}\setminus Z_{i_{0}}||W_{i_{0}}^{\prime}|-(3/4)\varepsilon n.$
Since $e_{G}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})\leq e_{T}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})+\sum_{j\neq{i_{0}}}e_{G}(Z_{j}^{i_{0}})$ and $\alpha_{i_{0}}=|\tilde{Z}_{i_{0}}\setminus Z_{i_{0}}||W_{i}^{\prime}|+e_{T}(W_{i_{0}}^{\prime})+e_{T}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})$, we have

### v2

@@ 3. $(K3)$
Suppose to the contrary, there is a vertex $v$ in $\tilde{U}_{1}\setminus(Z_{1}^{1}\cup W_{1}^{\prime})$ or $W_{1}^{\prime}\setminus V_{q_{1}}$, namely, $v\in Z_{i}^{1}$ for some $2\leq i\leq r$ or $v\in W_{1}^{\prime}\setminus V_{q_{1}}$. Suppose $v\in V_{l}$. Then $l\neq q_{1}$. Moreover, if $i$ is defined, then $W^{\prime}_{1}\cap V_{q_{1}}\subseteq V\setminus(\tilde{Z}_{i}\cup V_{l})$; otherwise, $W^{\prime}_{1}\cap V_{q_{1}}\subseteq V\setminus V_{l}$. By (Deg), we have $\overline{d}(v,W_{1}^{\prime}\cap V_{q_{1}})\leq 2\varepsilon n$. Let $N:=W_{1}^{\prime}\cap V_{q_{1}}\cap N(v)$. We have $|(W_{1}^{\prime}\cap V_{q_{1}})\setminus N|\leq 2\varepsilon N$. Since $|W_{1}^{\prime}\setminus V_{q_{1}}|\leq|W_{1}^{\prime}\setminus W_{1}|\leq 2\gamma n$, it follows that $|W^{\prime}_{1}\setminus N|\leq 2\varepsilon n+2\gamma n\leq 3\varepsilon n$.
Recall (4.13), $|W_{1}|>3t\varepsilon n$. By $(K3)$ (if $v\in\tilde{Z}_{1}\setminus Z_{1}$) or $(K1)$ (if $v\in W_{1}^{\prime}\setminus V_{q_{1}}$), we know that $G[Z_{1}^{1},N]$ contains no $K_{t-1,t}$ with the part of size $t$ in $N$. This implies that $e_{G}(Z_{1}^{1},N)=O(n^{2-1/(t-1)})$. Furthermore, by $(P3)$ and $(K1)$, $G[\tilde{Z}_{1}\setminus Z_{1}^{1},Z_{1}^{1}]$ is a $K_{t,t}$-free bipartite graph with one part of size at most $(r-1)\sqrt{\gamma}n$ and the other part of size at most $n$. Thus, $e_{G}(\tilde{Z}_{1}\setminus Z_{1}^{1},Z_{1}^{1})\leq C(r-1)\sqrt{\gamma}n^{2-1/t}$. By the similar arguments, we have $e_{G}(W_{1}^{\prime}\setminus N,Z_{1}^{1})\leq C(3\varepsilon n)n^{1-1/t}$.
Putting these bounds together (and note that $Z_{1}^{1}$ is an independent set, we get
$f_{1}$ $=e_{G}(Z_{1}^{1},N)+e_{G}(\tilde{Z}_{1}\setminus Z_{1}^{1},Z_{1}^{1})+e_{G}(W_{1}^{\prime}\setminus N,Z_{1}^{1})$
$=O(n^{2-1/(t-1)})+O(\sqrt{\gamma}n^{2-1/t})+O(\varepsilon n^{2-1/t}).$
By ($E1$), this contradicts (4.14). ∎
@@
$\alpha_{1}=\beta_{1}=0,\quad\text{and}\quad e_{G}(\tilde{U}_{1})=f_{1}\leq z_{t}(|Z_{1}^{1}|,|W^{\prime}_{1}|).$ (4.15)
Therefore, if $|W_{1}^{\prime}|\leq(1-\gamma)n$, then we have $f_{1}\leq z_{t}(n,|W_{1}^{\prime}|)\leq z_{t}(n,n)-\delta n^{2-1/t}$ for some $\delta>0$ by ($E2$). This contradicts (4.14). So we obtain
$\text{if }a=1,\text{ then }|W_{1}^{\prime}|\geq(1-\gamma)n\quad(\text{and thus}\quad 1\in L).$ (4.16)
@@
    If there is a copy of $K_{1,t}$ in $\tilde{U}_{i}\setminus(Z_{1}\cup V_{q_{1}})$, then there exists $l\in[r]\setminus\{i\}$ such that the vertex set of $K_{1,t}$ intersects both $V_{q_{l}}$ and $Z_{l}$.
@@
    Both $\tilde{Z}_{i}\setminus Z_{1}$ and $Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})$ are $K_{1,t}$-free.
@@
First assume that $N$ is $K_{t-1,t}$-free and thus $e_{G}(N)=O(n^{2-1/(t-1)})$. Note that, since $|\tilde{U}_{1}\setminus N|\leq(t+1)2\varepsilon n$, the edges in $\tilde{U}_{1}$ incident to $\tilde{U}_{1}\setminus N$ can be split into two bipartite $K_{t,t}$-free graphs each with one part of size at most $(t+1)2\varepsilon n$ and the other part of size at most $n$. Thus, the number of such edges is $O(\varepsilon n^{2-1/t})$. This gives $f_{1}=O(n^{2-1/(t-1)})+O(\varepsilon n^{2-1/t})$, contradicting (4.12).
@@
We thus assume $N$ contains a copy of $K_{t-1,t}$. Together with $B$, they form a copy of $K_{3}{(t)}$ in $G[\tilde{U}_{1}\cup\tilde{U}_{i}]$ and we denote its vertex set by $B^{\prime}$. By Claim 4.5, there exists $l\notin\{1,i\}$ such that $B^{\prime}$ intersects $V_{q_{l}}$ and $Z_{l}$. By Claim 4.8, $\tilde{U}_{1}\cap U_{l}=\emptyset$, so $B^{\prime}\cap Z_{l}=B\cap Z_{l}$ and $B$ indeed intersects $Z_{l}$. Since $\tilde{U}_{i}\cap Z_{l}\supseteq B\cap Z_{l}\neq\emptyset$, we infer that $|W_{l}|\geq(1-2\varepsilon)n$ from $(Q3)$, which implies that $q_{l}\neq q_{1}$ because of (4.13). It follows that $W_{1}\cap V_{q_{l}}=\emptyset$ and thus $B\cap V_{q_{l}}=B^{\prime}\cap V_{q_{l}}\neq\emptyset$, as desired.
@@
For Part (2), let $A_{i}:=Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})$ and $B$ be the vertex set of a copy of $K_{1,t}$ in $\tilde{Z}_{i}\setminus Z_{1}$ or in $A_{i}$. Then, by the first part of the claim, there exists $l\in[r]\setminus\{i\}$ such that $B$ intersects $V_{q_{l}}$ and $Z_{l}$. This is impossible if $B\subseteq A_{i}$ because $A_{i}\cap Z_{l}=\emptyset$ for any $l\notin\{1,i\}$, and also impossible if $B\subseteq\tilde{Z}_{i}\setminus Z_{1}$ because in which case $B\cap W=\emptyset$ and thus $B\cap V_{q_{l}}=\emptyset$ for any $l\notin\{1,i\}$. ∎
The following claim shows a clean structure for the $\tilde{U}_{i}$ such that $W_{i}^{\prime}$ is not too small.
@@
Suppose instead, for some $i_{0}\in[2,r]$ with $|W_{i_{0}}|\geq 2\varepsilon n$, there exists $v\in\tilde{U}_{i_{0}}\setminus(U_{i_{0}}\cup V_{q_{i_{0}}})$. By $(P4)$ and the fact that $v\in\tilde{U}_{i_{0}}\setminus U_{i_{0}}$, we infer that $d(v,Z_{j})\geq\varepsilon n$ for all $j\neq{i_{0}}$. Then, by Claim 4.7, we have $d(v,U_{i_{0}})<\varepsilon n$. Consequently, $d(v,W_{i_{0}}^{\prime}\cap W_{i_{0}})<\varepsilon n$, namely, $v$ has at least $2\varepsilon n-2\gamma n-\varepsilon n\geq(1/2)\varepsilon n$ non-neighbors in $W_{i_{0}}^{\prime}\cap W_{i_{0}}$ (in $G$). Note that $v$ is adjacent to all the vertices of $W_{i_{0}}^{\prime}\cap W_{i_{0}}$ in $T$. Since $G[W^{\prime}_{i_{0}}]\subseteq T[W^{\prime}_{i_{0}}]$, we infer that
$e_{G}(W_{i_{0}}^{\prime})+e_{G}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}},W_{i_{0}}^{\prime})\leq e_{T}(W_{i_{0}}^{\prime})+|\tilde{Z}_{i_{0}}\setminus Z_{i_{0}}||W_{i_{0}}^{\prime}|-(1/2)\varepsilon n.$
Since $e_{G}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})\leq e_{T}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})$ and $\alpha_{i_{0}}=|\tilde{Z}_{i_{0}}\setminus Z_{i_{0}}||W_{i}^{\prime}|+e_{T}(W_{i_{0}}^{\prime})+e_{T}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})$, we have